OpenStudy (anonymous):

(MEDAL REWARDED!) 5 1/3 Divided By 2/3

OpenStudy (anonymous):

\[5\tfrac{1}{3}=\frac{16}{3}\] is a start then dividing by \(\frac{2}{3}\) is the same as multiplying by \(\frac{3}{2}\) so compute \[\frac{16}{3}\times \frac{3}{2}\] cancel away

OpenStudy (anonymous):

Thanks